OpenStudy (anonymous):

does this series converge or diverge? 2n+1/n3^n explain

OpenStudy (anonymous):

it converges... using the divergence test: \[\lim_{n \rightarrow \infty}2n+1/n3^n=\lim_{n \rightarrow \infty}2n/n3^n+1/n3^n=\lim_{n \rightarrow \infty}2/3^n+1/n3^n=0+0\] =0, so the series converges

OpenStudy (experimentx):

not quite sure if it is right to conclude series converges using divergence test. http://www.math.oregonstate.edu/home/programs/undergrad/CalculusQuestStudyGuides/SandS/SeriesTests/divergence.html

OpenStudy (anonymous):

I suppose anonymous is correct but when the limit is zero using that test it doesn't necessarily mean it converges. Does it? Because the harmonic sequence limit is zero but it's a divergent series.

OpenStudy (experimentx):

i was thinking the same. but you can use comparison test. it's quite simplified above.

OpenStudy (anonymous):

@experimentX is very correct. The divergence test is only useful for testing the divergence of a series.

OpenStudy (anonymous):

or if you want to be absolutely sure, I will .... so: 2n+1/n3^n=2n/n3^n+1/n3^n=2/3^n+1/n3^n 2/3^n will converge since its a geometric series with ratio 1/3 for 1/n3^n, use ratio test so \[|(1/(n+1)3^{n+1})(n3^n)|=\lim_{n \rightarrow \infty}|n/3(n+1)|=\lim_{n \rightarrow \infty}|1/3|=1/3\] so that series is convergent, and since our series is a sum of two convergent series, then our series is also convergent